Balance the following in basic solution.

(a)Mn(OH)2(s) + O2(g) → MnO2(s)
Oxidation reaction
Reduction reaction
Net reaction

(b)In(s) + Cr2O72−(aq) → In(OH)3(s) + Cr(OH)4−(aq)
Oxidation reaction
Reduction reaction
Net reaction

(c)SO32−(aq) + Cu(OH)2(s) → SO42−(aq) + CuOH(s)
Oxidation reaction
Reduction reaction
Net reaction

(d)NO3−(aq) + H2(g) → NO(g)
Oxidation reaction
Reduction reaction
Net reaction
1
Expert's answer
2020-04-29T13:30:24-0400

(a) Mn(OH)2(s) + O2(g) → MnO2(s) 

Oxidation reaction

2Mn(OH)2 + 4OH- → 2MnO2 + 4H2O + 4e-

Reduction reaction

O2 + 4e- + 4H2O → 2H2O + 4OH- 

Net reaction

2Mn(OH)2 + O2 → 2MnO2 + 2H2O


(b) In(s) + Cr2O72−(aq) → In(OH)3(s) + Cr(OH)4(aq)

Oxidation reaction

2In + 6OH- = 2In(OH)3 + 6e-

Reduction reaction

Cr2O72- + 6e- + 7H2O = 2Cr(OH)4- + 6OH-

Net reaction

2In + Cr2O72- + 7H2O → 2In(OH)3 + 2Cr(OH)4-


(c) SO32−(aq) + Cu(OH)2(s) → SO42−(aq) + CuOH(s)

Oxidation reaction

SO32- + 2OH- = SO42- + 2e- + H2O

Reduction reaction

2Cu(OH)2 + 2e- = 2CuOH + 2OH-

Net reaction

SO32- + 2Cu(OH)2 → SO42- + 2CuOH + H2O


(d) NO3(aq) + H2(g) → NO(g)

Oxidation reaction

3H2 + 6OH- = 6H2O + 6e-

Reduction reaction

2NO3- + 6e- + 4H2O = 2NO + 8OH-

Net reaction

3H2 + 2NO3- = 2NO + 2H2O + 2OH-